OpenStudy (lposton):

What is the solution to -2(3x - 9) + 5x = -10? -28 -8 8 28

OpenStudy (idealist10):

-2(3x-9)+5x=-10 -6x+18+5x=-10 -6x+5x=-10-18 -x=-28 x=28 Answer: 28

OpenStudy (idealist10):

You just combine like terms in order to find the unknown variable x.

OpenStudy (lposton):

thank u

OpenStudy (idealist10):

You're welcome!